OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth if necessary. x^2 + 15 =0 A. {3.9} B. {3.9, -3.9} C. {7.5, -7.5} D. no real solution

OpenStudy (anonymous):

x^2=-15 the square root of a negative number has no real solutions

OpenStudy (anonymous):

D

OpenStudy (anonymous):

x^2 + 15 =0 x^2 = -15 no real solution

OpenStudy (anonymous):

 part in circle means you do not have a real solution.